A stone of weight 50N is completely immersed in a fluid. The weight of the stone shown on the spring balance is 48N. Calculate the volume of the stone

Given the density of fluid is 2100kg/m3


1
Expert's answer
2021-03-21T11:19:36-0400

"F_{G0}-F_G-F_B=0\to50-48=\rho Vg\to"


"V=\frac{2}{\rho g}=\frac{2}{2100\cdot9.8}=9.7\cdot10^{-5}\ (m^3)=97\ (cm^3)" . Answer